A 600-kg car is going around a curve with a radius of 120 m that is banked at an angle of 20° with a speed of 24.5 m/s. What is the minimum coefficient of static friction required for the car not to skid?


1
Expert's answer
2021-11-07T19:27:19-0500

"F_n \to" normal reaction Force

"F_F \to" frictional Force

"Mg \to" weight

"F_c \to" centrifugal Force

"\begin{cases}
 F_N -Mgcos20^o - F_Csin20^o = 0 \to F_N = Mgcos20^o + F_Csin20^o \\ 
 F_Ccos20^o - F_F-Mgsin20^o=0 
\end{cases}"

"F_Ccos20^o - F_F-Mgsin20^o=0"

"F_F=F_Ccos20^o -Mgsin20^o"

"\mu F_N=" "F_Ccos20^o -Mgsin20^o"

"\mu F_N=\frac{MV^2}{R}cos20^o-Mgsin20^o"

"\mu = \frac{1}{F_N}\lbrack\frac{MV^2}{R}cos20^o-Mgsin20^o\rbrack"

"\mu = \frac{1}{Mgcos20^o + F_Csin20^o}\lbrack\frac{MV^2}{R}cos20^o-Mgsin20^o\rbrack"

"\mu = 0.1235"